OpenStudy (anonymous):

Is my work correct...? 2x^2+16x+24 Factor the expression. I got (x+6)(x+2) because you divide everything by 2 and factor, right????

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

2(x+6)(x+2)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

2x^2+16x+24=2(x^2+8x+12=2(x+2)(x+6)

OpenStudy (anonymous):

oh ok